One-D Acceleration Histogram Points:4 A car is initially at rest on a straight road. The histogram below shows the cars acceleration along that road as a function of time. -1 -2 3 8123 45 67 8 9 18 11 12 13 14 Tine (s) Calculate the speed of the car at t 4 s. 12m/s Submit Answer Incorrect. Tries 4/7 Previous Tries Calculate the distance traveled during the first 6 s. Submit Answer Tries 0/7 Calculate the distance traveled from t-10 s to t 14 s. Submit Answer Tries 0/7 Calculate the cars average speed from t = 5 s to t=9 s Submit Answer Tries 0/7
Answer #1

a] here speed = area under a-t graph

= 3*[4-1]

= 9 m/s

b] speed at the end of t=6 s is 9-1*1 = 8 m/s

so distance travelled = 0.5*3*(4-1)^2 + 9*1 + 9*(6-5) - 0.5*1*(6-5)^2

= 31 m

c] speed at t= 10s, v = 9-1+4-2 = 10 m/s

distance travelled in given time = 10* [14-10]

= 40 m

d] distance covered =  9*(6-5) - 0.5*1*(6-5)^2 + 8*(8-6)+0.5*2*(8-6)^2 + 12*(9-8) -0.5*1*(9-8)^2

= 40 m

aaverage speed = distance/time = 40m/4s

= 10 m/s answer